OpenStudy (anonymous):

Assuming x > 0, what is the simplified fraction form for 15x times the square root of 3, all divided by the square root of 75 squared? Fan and Medal will be awarded!

OpenStudy (anonymous):

So the equation looks like \[\frac{15x \times \sqrt{3}}{ \sqrt{75^{2}} }\]

OpenStudy (anonymous):

I will award a medal and fan!

OpenStudy (matt101):

Well you can see the denominator becomes 75 right away. 15 and 75 reduce so you get (x√3)/5, and that's as far as you can go.

OpenStudy (anonymous):

\[\frac{ 15x \sqrt{3} }{ 75 }\] \[\frac{ x \sqrt{3} }{ 5 }\]

OpenStudy (anonymous):

Thanks guys(: